# secReferences in the Online Appendix

### v5

@@ 6.2 NGO Success and Reputation
where “$\text{Success}_{i,n,y}$” is defined as in (4)—namely, either a successful vote or a withdrawn proposal. On the right-hand side, “$\text{Campaign on Same Cause}_{i,n,y}$” is one if there has been a campaign by NGO $g$ against firm $n$ on cause $i$ in the ten months preceding the AGM and zero otherwise, and “$\text{Experience}_{g,y}$” indicates the number of campaigns NGO $g$ had launched in the past. We control for all confounding variations at the month, firm-by-year, NGO-by-year, and cause-by-year through fixed effects. Standard errors are clustered by firm.
@@
Main Variables
Campaign on Same Topic 0.008 0.076 0.032 0.070
@@
Experience 0.018
Campaign on Same Topic $\times$ Experience 0.026
AGM Experience (0/1) -0.042 -0.041
@@
Campaign on Same Topic $\times$ AGM Experience (0/1) 0.169∗∗ 0.150∗∗
@@
Non-AGM Experience (0/1) 0.005
Campaign on Same Topic $\times$ Non-AGM Experience (0/1) 0.065
AGM Experience -0.018
Squared AGM Experience 0.002∗∗
Campaign on Same Topic $\times$ AGM Experience 0.074∗∗
Campaign on Same Topic $\times$ Squared AGM Experience -0.005∗∗∗
@@
Firm x Year
Topic x Year
@@
Month
Fit statistics
Observations 38,613 38,613 38,613 38,613
R2 0.321 0.321 0.321 0.323
*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
@@
  • Note: This table reports the estimated coefficients from (6) in Column 1. In Column 1, Experience is capped at 5 to avoid outliers–as 85% of NGOs never reach that level of experience. In Column 2, we also consider the number of campaigns the NGO has launched in the past, which were not on the firm’s AGM day (Non-AGM Experience), and its interaction with Campaign on Same Topic. In Column 3 consider the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square (Squared AGM Experience), and their interaction with Campaign on Same Topic. All regressions also include firm-by-year, NGO-by-year, topic-by-year, and month fixed effects. Clustered (firm) standard errors are reported in parentheses.
Column (1) of Table 1 reports the estimated $\{\beta_{1},\beta_{2},\beta_{3}\}$. Campaigning on the same cause as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero. Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a cause being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
Column (2) unpacks Experienceg,y by considering the timing of an NGO’s past campaigns. We replace “$\text{Experience}_{g,y}$” for the variable “$\text{AGM Experience}_{g,y}$,” which is 1 if NGO $g$ launched at least one campaign on the day of the AGM in past years and zero otherwise. The interaction coefficient is significant and economically important. While a campaign on the same cause as the shareholder proposal increases its success at the AGM by 8pp (26%), the effect increases by 14pp (45%) when that NGO has a history of targeting AGMs.
[…]
Finally, as in signaling games where there are decreasing returns to signal one’s type if the agent has already signaled the type in the past, we estimate decreasing returns from launching campaigns on AGM dates. Column (4) replaces “$\text{Experience}_{g,y}$” for the cumulative count of the campaigns when NGO $g$ targeted an AGM date (“$\text{AGM Experience}_{g,y}$”) and its squared value: the last two rows of Column (4) find that the coefficients of the interactions between those variables and “$\text{Campaign on Same Cause}_{i,n,g,y}$” imply a concave relationship between past AGM-targeting and proposals’ success. Therefore, the support of NGO’s to a shareholder proposal display diminishing returns. [^13] Note that the combined effect of AGM Experience and AGM Experience Squared never yields a negative probability of proposal success—even at the highest observed levels of AGM Experience—suggesting our reputation proxy does not exhibit negative returns.
This finding supports the results in Appendix Figure B4, discussed in Section 4, which demonstrate that AGM-targeting increases the probability that target receives negative media coverage only if the NGO has little campaign experience (i.e., having run less than four campaigns). Later campaigns do not generate additional awareness on AGM dates compared to other dates. In fact, “mature NGOs” seem to be better off campaigning one or two months earlier, indicating diminishing returns from campaigning on AGM dates, even in terms of visibility.
We conclude that,
@@
Targeting AGMs helps an NGO driving support for shareholder proposals in future periods, but displays decreasing returns.
@@ 7 Discussion and Conclusion
We first show that NGOs tend to launch their campaigns on the AGM date, as media coverage is higher during AGMs (Result 1). By doing so, NGOs elicit reactions from key stakeholders, such as shareholders and consumers (Result 2). The immediate goal of these campaigns is to influence a firm’s strategy by damaging its reputation and voicing stakeholder dissatisfaction.
@@
However, campaigning before the AGM offers direct advantages. Specifically, when an NGO targets a firm on an issue for which a shareholder proposal has already been filed, it boosts the proposal’s chance of acceptance—often leading to its withdrawal from the ballot—and raises average shareholder support by 8% (Result 3). Given that proposals typically attract low support, early NGO interventions can effectively discipline managers (Levit and Malenko, 2011). Importantly, all these analyses control for NGO-by-year fixed effects and cause-by-year fixed effects, accounting for any time-varying confounding factors at the NGO level—such as experience or reputation—or at the cause level—such as specific news events—thus ensuring that globally renown NGOs (e.g., PETA) do not drive these results. Furthermore, we show that selection issues are not a central concern: Panel d of Appendix Table A1 shows target‐firm characteristics are invariant to an NGO’s past campaign experience, and Appendix Table A2 confirms NGOs campaigning at AGMs do not systematically target different firms than those campaigning at other times.
@@
Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time. NGOs with limited activism histories tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4). One interpretation is that as NGOs gain more experience with campaigning, they accumulate media exposure, reducing the pressure to campaign specifically on the AGM date.
@@
Therefore, repeatedly targeting AGMs may yield diminishing returns (Result 5). First, an NGO may have already signaled its commitment to a cause, making subsequent signals weaker and reducing the effectiveness of AGM-date campaigning as a signaling device. This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following ”reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future. [^14] In these models, incumbent firms may incur short-term losses by setting predatory prices upon entry of a rival to deter future entrants (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
@@
Second, stakeholders may come to expect more constructive approaches and proposals from established NGOs, leading NGOs to support earlier shareholder proposals or other unobserved tactics. Consistent with this, mature NGOs—those with multiple prior campaigns—do not inflict greater reputation harm by campaigning on the AGM date, whereas they do so in their initial campaigns (compare the blue dots and red triangles in Appendix Figure B4). In fact, experienced NGOs generate the same or even more negative media coverage for a firm by launching campaigns one or two months before the AGM.
@@
In conclusion, NGOs influence corporate behavior, yet their incentives and strategies remain understudied. This paper analyzes NGO campaigns, leveraging the exogeneity of the corporate calendar to uncover their trade-offs. Similar approaches have identified stakeholder preferences, showing that AGMs are more than voting venues—they also serve as forums for firm-stakeholder negotiations, often unseen by researchers but inferred from underlying incentives (Fioretti et al., 2023).
@@ Appendix A Omitted Tables
      – E campaigns 1.51 0 1 1 5
      – S campaigns 1.70 0 1 1 6
@@
E score 5.66 4 5.58 7.3 2.22
S score 4.44 3.30 4.70 5.60 1.58
E campaign 0.47 0 0 1 0.50
S campaign 0.53 0 1 1 0.50
@@
High Experience Low Experience Test difference
@@
E score 5.64 0.12 5.72 0.06 0.58
S score 4.44 0.09 4.45 0.04 0.89
E campaign 0.46 0.03 0.5 0.01 0.15
S campaign 0.54 0.03 0.49 0.01 0.12
Panel d: OLS Regression of $y_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
@@
E score 0.004 (0.009) 0.041 (0.041) 1,338
S score -0.0006 (0.006) 0.019 (0.032) 1,338
E campaign 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
S campaign -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
  • Note: This table reports summary statistics on NGO campaigns (Panel a, NGO campaign data from Sigwatch and AGM dates from ISS), firms targeted by the campaigns (Panel b, accounting data from Compustat, ESG scores from MSCI, and news count based on RavenPack data), as well as an analysis of the relationship between NGO experience and targeted firms’ characteristics (Panels c and d). In Panel c, the Market Cap, Total Assets, and Sales deciles are based on the universe of US firms in Compustat on the quarter of the campaign. We use deciles instead of dollar values because highly ranked firms saw a large expansion in market cap and assets over the last years. Ranking measures like deciles avoid such a mechanical correlation with NGO Experience, which, being defined as the number of campaigns launched by NGO g in previous years, also grows over time. In Panel d, the fixed-effects are: NGO×$\times$×Year, Firm×$\times$×Year and Topic×$\times$×Year, except for the last two rows where Topic×$\times$×Year fixed effects are not included because of perfect collinearity with the dependent variables. Standard errors are clustered by firm.
Table A2: Campaign timing and characteristics of target firms
Characteristics of Target Firm
Market Cap Assets Sales Yearly Environmental Social Campaign
Dependent Variables: (decile) (decile) (decile) News Score Score Sentiment
Model: (1) (2) (3) (4) (5) (6) (7)
Variables
Campaign on AGM day (0/1) 0.469 0.495 0.374 83.2 0.469 -0.522 0.022
(0.628) (0.695) (0.550) (185.7) (0.919) (0.715) (0.383)
Fixed-effects
Industry-Year
NGO-Year
Topic-Year
Fit statistics
Observations 1,338 1,338 1,337 1,299 1,338 1,338 1,338
R2 0.93364 0.92669 0.90388 0.99232 0.95420 0.92167 0.84525
*–p$<0.1$; **–p$<0.05$; ***–p$<0.01$.
  • Note: This table reports the estimated coefficients from regressions of different target firm characteristics on a dummy indicating whether the campaign took place on the target firm’s AGM day. All regressions also include Industry-by-Year, NGO-by-Year, and Topic-by-Year fixed effects. The Market Cap, Assets, and Sales deciles are based on the entire universe of US firms in Compustat in the quarter of the campaign. Yearly News refers to the four-quarter moving average of news mentioning the target firm. Accounting data from Compustat. NGO campaign (including Sentiment) data from Sigwatch. News data from RavenPack. Environmental and Social scores are from MSCI. Standard errors are clustered by firm.
@@ Appendix B Omitted Figures
Figure B1: Count of E&S shareholder proposals and their success
Note: This figure plots the number of E&S proposals at Russell 3,000 firms from 2006 to 2021 (solid line) and the number of successful E&S proposals (blue area) in the same period. A successful proposal is defined as a proposal that is either withdrawn before the AGM, or that goes to vote and passes. Data from ISS. Based on 5,937 shareholder proposals filed at 979 firms.
@@
Figure B3: Testing NGO campaign timing around the AGM date
Note: This figure plots the density test for the occurrence of NGO campaigns around firms’ AGM, following Cattaneo et al. (2020). The test is based on 521 NGO campaigns launched within 50 days of the firm’s closest AGM. The NGO data come from Sigwatch; the AGM dates are from ISS.
Figure B4: Media exposure of NGO campaigns around an AGM
(a) Main specification
(b) Controlling also for current E&S news
Note: The figure presents estimates from (1). The dependent variable is an indicator that a firm receives at least one negative E&S news in a week. Panel (b) also controls for the current count of E&S news to account for firm-week specific events that increase news coverage (not necessarily negative coverage). Standard errors are clustered at the firm level.
Figure B5: Examples of positive and negative E&S news
(a) Example of a positive E or S news
(b) Example of a negative E or S news
Notes: This figure displays examples of a positive (top) and negative (bottom) E&S news. News data from RavenPack.
Figure B6: Media sentiment of NGO campaigns around an AGM
Note:The figure presents estimates from (1). The dependent variable is the average sentiment of E&S news received by firm in a week. Standard errors are clustered at the firm level.
@@
Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead. The regression also includes month-by-year-by-industry-by-topic fixed effects. The regression is based on 1,227 NGO campaigns. Standard errors are clustered at the firm level.
Figure B8: NGO campaigns and consumer backlash
(a) NGO Campaigns on AGM
@@
(c) Estimated Effect of Campaigning on AGM
Notes: Panels (a) and (b) plot the estimated coefficients of the regression of sales over (lagged) total assets, on a set of time dummies interacted with a dummy indicating the occurrence of an NGO campaign targeting firm $n$ from (3). Panel(c) plots the estimated $\hat{\beta}$ from
$\frac{\text{Sales}_{n,q}}{\text{Assets}_{n,q-1}}$ $=\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\beta_{\tau}\cdot\text{NGO Campaign on AGM}_{n}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)$
$+\sum^{\tau=5}_{\begin{subarray}{c}\tau=-4\\ \tau\neq-1\end{subarray}}\left(\tilde{\beta}_{\tau}\cdot\mathbf{1}{\{\text{distance}=\tau\}}\right)+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$
which interacts the time dummy, the occurrence of a campaign, and a dummy indicating that the campaign happened on AGM day. The regressions also include firm and quarter-by-industry fixed effects. The reference period is -1 quarter, consistent with the other analyses. The number of observations is 10,188, 351,351, and 361,539, respectively. Standard errors are clustered at the firm level.
@@
$\Delta_{4}\text{E\&S Score}_{n,q}=\sum_{l=-6}^{5}\text{Campaign}_{n,q,l}+\textbf{X}_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
where $\Delta_{4}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter $q$ and quarter $q+4$, and $\sum_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM. The regression also controls for firm size (logged total assets), current average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$). ESG scores are from MSCI, NGO campaigns data from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS. The regression has 47,082 observations. Standard errors are clustered at the firm level.
@@
Appendix C Matching Proposals And NGO Campaigns

### v6

@@ 6.3 NGO Reputation and AGM Influence
where $\text{Success}_{i,n,y}$” indicates whether a proposal is withdrawn or passes a vote, as in equation (4). The key variable $\text{Campaign on Same Cause}_{i,n,y}$” equals one if NGO $g$ campaigned against firm $n$ on cause $i$ in the ten months preceding the AGM. $\text{Experience}_{g,y-1}$” counts NGO $g$’s past campaigns. We absorb firm-by-year, NGO-by-year, cause-by-year, and month variation through fixed effects, and cluster standard errors at the firm level.
@@
Campaign on Same Topic -0.0003 -0.013 -0.005 -0.001
@@
Experience -0.034
Campaign on Same Topic $\times$ Experience 0.078∗∗∗
Experience square 0.008∗∗
Campaign on Same Topic $\times$ Experience square -0.015∗∗
Non-AGM Experience 0.0008
Campaign on Same Topic $\times$ Non-AGM Experience -0.011
AGM Experience 0.005 -0.034
@@
Campaign on Same Topic $\times$ AGM Experience 0.048 0.091∗∗∗
@@
AGM Experience square 0.010∗∗
Campaign on Same Topic $\times$ AGM Experience square -0.013∗∗
Campaign on Same Topic $\times$ Google Search Popularity 0.071
Campaign on Same Topic $\times$ Google Search Popularity square -0.021
@@
CUSIP x Year
@@
Topic x Month-Year
Observations 36,543 35,778 36,543 33,142
R2 0.61894 0.61698 0.61995 0.62078
*–p<0.1<0.1; **–p<0.05<0.05; ***–p<0.01<0.01.
@@
  • Note: This table reports the estimated coefficients from (7) in Column 1. In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (Non-AGM Experience) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), and their interaction with Campaign on Same Topic. In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square, and their interaction with Campaign on Same Topic. In Column 4, we replace Experience from model 7 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign. All regressions also include firm-by-year, NGO-by-year, topic-by-month-by-year fixed effects. Clustered (firm) standard errors are reported in parentheses.
Column (1) of Table 1 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$). What matters is the interaction: proposals are more likely to succeed when the campaigning NGO has accumulated experience ($\beta_{3}>0$). A one-standard-deviation increase in past campaigns on the same cause raises proposal success by about 8% on average.
Column (2) distinguishes between campaigns held on AGM dates and those held at other times. Only the former contribute to proposal success. In fact, one-standard-deviation more AGM campaigns increases the probability of success by roughly 4.5 percentage points (about 15%), while non-AGM campaigns have no measurable effect.
Column (3) tests for diminishing returns by including squared terms. The results confirm a concave pattern: additional AGM campaigns continue to improve proposal success, but at a decreasing rate. Importantly, the combined effect remains non-negative even for the most active NGOs. [^8] The marginal effect of AGM Experience and its square never turns negative, so our proxy for reputation does not imply adverse returns.
We corroborate these findings with an alternative proxy: NGO visibility measured by Google Trends. Specifically, Column (4) replaces experience by Google Search Popularityg,y, which we define as the average search score of NGO $g$ relative to Greenpeace in year y1y-1.999The table drops the direct effect of this variable on the dependent variable, as it is absorbed by NGO-by-Year fixed effects. The interaction between this measure and campaigning on the same cause is positive, with a concave shape. Thus, both proxies—past AGM campaigns and Google visibility—suggest that reputation enhances the effectiveness of NGO campaigns, albeit with diminishing marginal returns.
These results also resonate with the media evidence in Section 4: AGM targeting increases negative media coverage only for relatively inexperienced NGOs. More seasoned NGOs gain little additional visibility from campaigning exactly on AGM dates and may be more effective when campaigning slightly earlier. These results support our view that:
@@
Targeting AGMs builds NGO reputation and increases their ability to raise the supported of related shareholder proposals, but with diminishing returns.
@@ 7 Discussion and Conclusion
We first show that NGOs tend to launch their campaigns on the AGM date, likely motivated by the higher media coverage surrounding these events (Result 1). By doing so, NGOs elicit reactions from key stakeholders, such as shareholders. However, the immediate purpose of these campaigns is primarily to gain visibility by voicing stakeholder concerns; they have no direct effect on firms’ current E&S scores, which improve only when campaigns are launched ahead of the AGM (Result 2).
@@
Hence, campaigning before the AGM offers direct advantages. When an NGO targets a firm on an issue for which a shareholder proposal has already been filed, it boosts the proposal’s chance of success—often leading to its withdrawal from the ballot—and raises average shareholder support by 8% (Result 3). Given the typically low baseline support for such proposals, early NGO interventions can effectively discipline managers who must contend with heightened shareholder pressure (Levit and Malenko, 2011). Importantly, all analyses include NGO-by-year and cause-by-year fixed effects, thereby accounting for time-varying confounders at both the NGO level—such as experience or reputation—and the cause level—such as specific news events. This ensures that globally renowned NGOs (e.g., PETA) do not drive the results. Furthermore, the analyses in Sections 4 and 6 show little evidence of selection on observables, alleviating concerns that NGOs disproportionately target firms that are easier to influence or more likely to generate visibility.
@@
Our results suggest that NGOs face a trade-off, which they navigate by adapting their strategies over time. NGOs with limited activism histories (or low visibility) tend to campaign on the AGM date, whereas more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4). One interpretation is that as NGOs gain more experience with campaigning, they accumulate media exposure, reducing the pressure to campaign specifically on the AGM date.
@@
Therefore, repeatedly targeting AGMs may yield diminishing returns (Result 5). First, an NGO may have already signaled its commitment to a cause, making subsequent signals weaker and reducing the effectiveness of AGM-date campaigning as a signaling device. This interpretation aligns with the view that AGM-date campaigns enhance an NGO’s “reputation.” For instance, following “reputation-for-toughness” models, the NGO may accept an opportunity cost—delaying support for a shareholder proposal—to launch a campaign at a time when it imposes the greatest costs on a firm, thereby building a reputation that can be leveraged with other targets in the future. [^10] In these models, incumbent firms may incur short-term losses by setting predatory prices upon entry of a rival to deter future entrants (Kreps and Wilson, 1982; Milgrom and Roberts, 1982).
@@
Second, stakeholders may come to expect more constructive approaches from established NGOs, such as supporting related shareholder proposals. Consistent with this hypothesis, AGM-date campaigns by mature NGOs—those with multiple prior campaigns—do not generate more negative media coverage than campaigns at other times. By contrast, the very same NGOs do so in their initial campaigns (compare the blue dots and red triangles in Appendix Figure B10). Hence, we observe not only a shift in strategy as NGOs mature (Figure 4) but also a change in their capacity to inflict reputational harm on firms.
@@
In conclusion, NGOs influence corporate behavior, yet their incentives and strategies remain understudied. We leverage the exogeneity of corporate events to uncover a trade-off between visibility and influence that NGOs face when deciding when to campaign, and show that how this trade-off is resolved determines which stakeholders are mobilized (e.g., shareholders today versus shareholders tomorrow). Hence, NGOs act as a coordination device, with distinct implications for the social and environmental impact of targeted firms.
@@ Appendix A Omitted Tables
     – Campaigns on Env. topic 1.51 0 1 1 5
     – Campaigns on Soc. topic 1.70 0 1 1 6
@@
Environmental score 5.66 4 5.58 7.3 2.22
Social score 4.44 3.30 4.70 5.60 1.58
Campaign on Env. topic 0.47 0 0 1 0.50
Campaign on Soc. topic 0.53 0 1 1 0.50
@@
NGO w/ AGM Experience NGO w/o AGM Experience Test difference
@@
Environmental score 5.64 0.12 5.72 0.06 0.58
Social score 4.44 0.09 4.45 0.04 0.89
Campaign on Env. topic 0.46 0.03 0.5 0.01 0.15
Campaign on Soc. topic 0.54 0.03 0.49 0.01 0.12
Panel d: OLS Regression of $z_{ny}=\beta_{0}+\beta_{1}\cdot\textit{NGO Experience}_{g(n)y}+\varepsilon_{ny}$
@@
Environmental score 0.004 (0.009) 0.041 (0.041) 1,338
Social score -0.0006 (0.006) 0.019 (0.032) 1,338
Campaign on Env. topic 0.006∗∗∗ (0.002) 0.021∗∗ (0.010) 1,338
Campaign on Soc. topic -0.006∗∗∗ (0.002) -0.018 (0.012) 1,338
[…]
@@ Appendix B Omitted Figures
Figure B2: Firms’ negative media coverage increases if NGOs campaign during AGMs
Note: The figure plots the coefficients $\beta_{\tau}$ from the following event study regression where we re-calendarize time around an AGM date so that the AGM month is month 0 in each year for each firm:
$News_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\!\left(\beta_{\tau}\cdot 1\!\left\{M_{n,w,y}=\tau\!\right\}\cdot NGO\,Campaign_{n,w,y}\!\right)$
$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
where $News_{n,w,y}$ is an indicator for whether firm $n$ in week $w$ of year $y$ experiences at least one negative E&S news event. The set of dummies $1\{M_{n,w,y}=\tau\}$ indicate the distance in months from the closest AGM. The function $M_{n,w,y}$ assigns week $w$ to each month; the indicator is 1 if the week falls in month $\tau$, where a month is made up of 4 weeks, and zero otherwise. We interact these event-study indicators, with the variable $NGO\,Campaign_{n,w,y}$, an indicator variable equal to 1 if firm n experienced an NGO campaign in week $w$ of year $y$, and 0 otherwise. The vector $X_{n,w,y}$ includes the direct effect of these variables. The reference month is $\tau=-4$ following SEC rules on shareholder proposals. Note that, because of the stability of AGM dates across years within firms (Appendix Figure B1), the distribution of months fixed effects is unaffected by the re-calendarization, so that months are not double counted across years. Standard errors are clustered by firm.
@@
Note: This figure plots the estimated coefficients of the regression of a dummy indicating the occurrence of a shareholder proposal on month $t$ at firm $f$, relative to an NGO campaign targeting firm $f$ on the same cause, $k$ months ahead. The regression also includes campaign and month-by-year-by-industry-by-topic fixed effects. The regression is based on 1,227 NGO campaigns. Standard errors are clustered at the firm level.
Figure B4: NGO campaign timing and consumer backlash
(a) NGO Campaigns on AGM dates
@@
Notes: The figure plots the estimated ${\beta}_{\tau}$ from (3). Panel (a) focuses on NGO campaigns on AGM dates, while Panel (b) focuses on campaigns on dates other than AGM dates. The reference period is -1 quarter, consistent with the other analyses. The number of observations is 239,818 in Panel (a) and 239,925 in Panel (b). Standard errors are clustered at the firm level.
@@
$_{5}\text{E\&S Score}_{n,q}=\sumop\displaylimits_{l=-6}^{5}\text{Campaign}_{n,q,l}+X_{n,q}\gamma^{\prime}+\alpha_{n}+\alpha_{s(n),q}+u_{n,q},$ (B1)
where $_{5}\text{E\&S Score}_{i,q}$ is the change in firm $n$’s average of its Environmental and Social scores between quarter q1q-1 and quarter $q+4$, and $\sumop\displaylimits_{l=-6}^{6}\text{Campaign}_{n,q,l}$ is a set of dummy variables indicating the occurrence of an NGO campaign against firm $n$ in quarter $q$, being $l$ months away from the closest AGM. The regression also controls ($X_{n,q}$) for firm size (logged total assets), lagged average of E and S scores, firm fixed effects ($\alpha_{n}$), and industry-by-quarter fixed effects ($\alpha_{s(n),q}$). ESG scores are from MSCI, NGO campaigns data from Sigwatch, firm characteristics from Compustat, and AGM dates from ISS. The regression has 47,082 observations. Standard errors are clustered at the firm level.
@@
Figure B10: Media coverage of NGO-target firms around their AGM
Note: To test if media exposure of NGO campaigns is highest during the AGM week, we aggregate negative E&S news by firm ($n$) and week-by-year ($w,y$) and perform the event study:
$\text{News}_{n,w,y}$ $=\sumop\displaylimits_{\begin{subarray}{c}\tau=-5\\ \tau\neq-4\end{subarray}}^{5}\left(\beta_{\tau}\cdot\mathbf{1}_{\{M_{n,w,y}=\tau\}}\cdot\text{NGO Campaign}_{n,w,y}\right)$
$\qquad+X_{n,w,y}\gamma^{\prime}+\alpha_{n,y}+\alpha_{w,y,s(n)}+u_{n,w,y},$
[…]
Appendix C Robustness of Google Trends Results
This section discusses the robustness of the results presented in Figure 2 to alternative specifications. In particular, we replicate the analysis using (i) a staggered difference-in-differences estimator (Sun and Abraham, 2021), which accounts for heterogeneous campaign timing across NGOs, and (ii) a distributed lags specification, which controls for the possibility of overlapping campaigns within the same year.
C.1 Staggered Difference-in-Differences Specification
We first replicate Figure 2, applying the staggered difference-in-differences approach to account for heterogeneous campaign timing across NGOs (Sun and Abraham, 2021).
Estimated coefficients $\beta_{\tau}^{AGM}$ are reported in Figure C1. The results are consistent with our baseline findings: AGM-timed campaigns (blue dots) generate sizeable visibility gains for early campaigns, raising Google search intensity by about 58%. By contrast, experienced NGOs see a more modest increase (33%) in search intensity, and only in the month of the campaign.
Figure C1: Staggered difference-in-differences specification
Note: The Figure plots the estimated $\beta^{AGM}$ coefficients from (1) using a staggered difference-in-differences estimator (Sun and Abraham, 2021). The dependent variable is the monthly Google Trends score for the NGO. Blue dots are estimated by restricting the sample to an NGO’s first three campaigns, while red triangles consider later campaigns. Standard errors are clustered at the campaign level.
C.2 Distributed Lags Specification
In the baseline event-study, if an NGO launches several campaigns in close succession, their effects may overlap in calendar time. For example, if one campaign takes place at $t=0$ and another at $t=2$, the months around $t=2$ capture the lingering impact of both events. This overlap can bias the estimated effects in the previous analyses. To address this, we estimate a distributed lags specification that explicitly includes leads and lags of both AGM and non-AGM campaigns:
$\text{Google Trends Score}_{g,m,y}$ $=\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-5\end{subarray}}\left(\beta_{\tau}^{AGM}\cdot\text{AGM Campaign}_{g,m+\tau}\right)$ (C1)
$\quad+\sumop\displaylimits^{\tau=5}_{\begin{subarray}{c}\tau=-5\end{subarray}}\left(\beta_{\tau}^{Non-AGM}\cdot\text{Non-AGM Campaign}_{g,m+\tau}\right)$
$\quad+\alpha_{m,y}+\alpha_{g,y}+u_{g,m,y},$
By doing so, the regression attributes observed variation in search activity in each month to the full sequence of campaigns occurring in that window, rather than assigning the entire effect to whichever campaign happens to be closest in event time. Put differently, the distributed lags approach disentangles the impact of overlapping campaigns by allowing each campaign to contribute separately to visibility over several months. This ensures that the estimated $\beta^{AGM}_{\tau}$ captures the marginal visibility effect of a single campaign, net of any other campaigns that may have occurred nearby.
Figure C2: Distributed lags specification
(a) First three campaigns
(b) Later campaigns
Note: The two plots present the estimated $\beta^{AGM}$ with blue circles and $\beta^{Non-AGM}$ with red triangles coefficients from (C1) on two different samples. The first panel subsets the data to each NGO’s first three campaigns, while the second panel examines later campaings. The dependent variable is the monthly Google Trends score for the NGO. Standard errors are clustered at the campaign level.
Figure C2 reports the estimated coefficients. When restricting attention to an NGO’s first three campaigns, the results confirm that AGM-timed campaigns lead to significant short-run increases in visibility (blue dots), while later campaigns generate little or no effect. Whether early or late, non-AGM campaings have a limited effect on an NGO’s visibility.
Overall, these robustness checks confirm that our main results are not driven by campaign timing heterogeneity across NGOs or by overlapping campaigns within a year.
Appendix D Matching Proposals And NGO Campaigns